three times the sum of a number n and 7 is 16

Answer by checkley79(3341)

3(N+7)=17
3N+21=17
3N=17-21
3N=-4
N=-4/3 ANS.
PROOF:
3(-4/3+7)=17
3(-4/3+21/3)=17
3*17/3=17
51/3=17
17=17